OpenStudy (anonymous):

Solve the quadratic equation. 48 - 8x2 = 0 Check your solution by substituting your answer back in the equation. A. ± √6 B. ± 2√2 C. ± (2√2)/√2 D. ± (4√3)/(2√2)

OpenStudy (anonymous):

The easiest way to solve something like this when they ask you to plug in the solution is to plug in the answers.

OpenStudy (anonymous):

48 - 8(sqrt6)^2 = 0 48 - 8(6) = 0 48 - 48 = 0 True So the correct answer is A.

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

No problem buddy. :)